Could you please tell me how to do these types of questions...
7x(squared)+9x+2
They are called "Algebraic Factorisation" Could you give me an example perhaps it is for yoar 10 mathematics.

Answer by ishkumargupta(1)

7x(squared)+9x+2
=7x(squared)+7x+2x+2 (write as 9x=7x+2x)
=7x(x+1)+2(x+1) 7x(squared)+7x = 7x(x+1) , 2x+2=2(x+1)
=(x+1)(7x+2) take x+1 common from 7x(x+1)+2(x+1)
Answer 7x(squared)+9x+2 =(x+1)(7x+2)